# A Mess of Memory System Benchmarking, Simulation and Application Profiling

### v1

@@ Abstract
The current Mess benchmark release covers all major CPU and GPU ISAs, x86, ARM, Power, RISC-V, and NVIDIA’s PTX. We also release as open source the ZSim, gem5 and OpenPiton Metro-MPI integrated with the Mess memory simulator for DDR4, DDR5, Optane, HBM2, HBM2E and CXL memory expanders. The Mess application profiling is already integrated into a suite of production HPC performance analysis tools.
@@ I Introduction
Our study argues that the memory system benchmarking, simulation and application profiling can and should be based on a unified view of memory system performance. We provide this view with the Memory stress (Mess) framework comprised of the Mess benchmark, memory simulator and application profiling tool (Figure 1). Mess benchmark (Section II) describes the memory system performance with a family of bandwidth–latency curves. The benchmark covers the full range of the memory traffic intensity, from the unloaded to fully-saturated memory system. It also considers numerous compositions of read and write operations, plotted with different shades of blue in Figure 1 (middle). The Mess benchmark is designed for holistic and detailed memory system characterization that is easily adaptive to different target platforms. The current benchmark release covers all major CPU and GPU ISAs, x86, ARM, Power, RISC-V, and NVIDIA’s Parallel Thread Execution (PTX), and it is applied to a number of actual hardware platforms and simulators.
@@
We deploy the Mess benchmark to characterize Intel, AMD, IBM, Fujitsu and Amazon servers as well as NVIDIA GPUs with DDR4, DDR5, HBM2 and HBM2E memory (Section II-D). We report and discuss a wide range of memory system behavior even for the hardware platforms based on the same standard. These differences are especially pronounced in the high-bandwidth areas which have the greatest impact on memory-intensive applications. To the best of our knowledge, we are the first ones to detect and analyze the effect of memory system over-saturation, a scenario in which further increase of the memory request rate causes the system performance drop.
@@
We also use the Mess benchmark to evaluate memory models of the widely-used hardware simulators: event-based ZSim [15], cycle-accurate gem5 [16] and RTL simulator OpenPiton Metro-MPI [32] (Section II-E). Unfortunately, all tested memory simulators, including well-established and trusted gem5 DDR models and cycle-accurate memory simulators [22, 24], poorly resemble the actual system performance. The simulators show an unrealistically low load-to-use latency (starting at 4 ns), high memory bandwidth (exceeding 1.8$\times$ the maximum theoretical one), and a simulation error of tens of percents for memory-intensive benchmarks, STREAM [5], LMbench [8] and Google Although it was not the initial design target, we also detected a case in which holistic and detailed Mess benchmarking diagnosed a CPU design error. In particular, Mess benchmarking of the OpenPiton measured an unexpected memory traffic, which led us to a bug in the coherency protocol generated by the OpenPiton framework.
@@
Apart from the memory system characterization, the Mess bandwidth–latency curves can be also used for the memory performance simulation (Section III). We develop and integrate the Mess memory simulator in the ZSim, gem5, and OpenPiton Metro-MPI simulators, enabling simulation of high-end memory systems based on DDR4, DDR5, Optane, HBM2, and HBM2E technologies, and Compute Express Link (CXL) [33]. The Mess integration is easy, based on the standard interfaces in the CPU simulators that enable external memory simulation [15, 16, 34]. The Mess simulator closely matches the actual memory systems performance. The simulation error for memory-intensive STREAM, LMbench and Google multichase is between 0.4% and 6%, which is significantly better than all other memory models we tested. Mess is also fast. For example, Mess integrated with ZSim introduces only 26% simulation time increment over the fixed-latency memory model, while the speed-up over the Ramulator and DRAMsim3 ranges between 13$\times$ and 15$\times$. Another advantage of the Mess simulator is that it removes the current lag between the emergence of memory technologies and development of the reliable simulation models. For example, Mess is the first memory simulator that models CXL memory expanders, enabling further research on these novel memory devices. The simulation is based on the bandwidth–latency curves obtained from Micron Technology’s SystemC hardware model (Section III-C).
Finally, the Mess memory bandwidth–latency curves enhance application profiling and performance analysis (Section IV). Mess application profiling determine positions of application execution time segments on the corresponding memory bandwidth–latency curves. The application memory stress can be combined with the overall application timeline analysis and can be linked to the source code. The Mess application profiling is already integrated into a suite of production HPC performance analysis tools [35].
@@
The inherent dependency between the used memory bandwidth and the access latency is by no means a new phenomenon, and the community has known about it at least for a couple of decades [36]. The Mess framework extends the previous work in three important aspects. Previous studies typically use a single bandwidth–latency memory curve to illustrate a general memory system behavior [36, 27, 12, 37, 38, 39]. The Mess benchmark is designed for holistic, precise and close-to-the-hardware memory system performance characterization. To reach this objective, the Mess benchmark is developed directly in the assembly, and the experiments are tailored to minimize and mitigate the impact of the system software. Mess benchmark performs a detailed characterization of the particular system under study. It detects and quantifies some aspects of the memory systems behavior not discussed in previous studies, such as the impact of the read and write memory traffic on performance, or discrepancies between different memory systems, both in actual platforms and hardware simulators. Second, the Mess framework tightly integrates the memory performance characterization into the memory simulation. The Mess simulator avoids complex memory system simulation and simply adjusts the rate of memory instructions (provided by the CPU simulator) to the actual memory performance. The accuracy of this simulation approach relies on the input memory performance characterization. This characterization therefore has to be holistic, detailed and specific to the system under study, closely matching the Mess benchmark design. Third, the framework closely couples the memory-related profiling of hardware platforms and applications. Similar to the Mess simulator, the Mess application profiling itself is uncomplicated, and its real value comes from the application analysis in the context of the memory system characteristic. The quality of the memory characterization therefore directly impacts the the quality of the overall analysis. For this reason, the analysis cannot be performed with generic bandwidth–latency approximations, but it requires detailed Mess-like memory performance description.
@@
The Mess benchmark is released as open source and it is ready to be used in x86, Power, ARM, and RISC-V CPUs and NVIDIA GPUs. The release also contains all bandwidth–latency measurements shown in the paper, including the CXL expander curves provided by Micron Technology. We also release as open source the ZSim, gem5 and OpenPiton Metro-MPI integrated with the Mess memory simulator supporting DDR4, DDR5, Optane, HBM2, HBM2E and CXL expanders. Also, the public releases of the production HPC performance analysis tools already include the Mess application profiling extension. The released tools are ready to be used by the community for better understanding of the current and exploration of future memory systems.
@@ II-A Mess benchmark: Memory bandwidth–latency curves
The Mess memory characterization comprises tens of bandwidth–latency curves, each corresponding to a specific ratio between the read and write memory traffic. The Mess benchmark kernels cover the whole range of memory operations, from 100%-loads to 100%-stores. The 100%-load kernel generates a 100%-read memory traffic, while the 100%-store kernel creates 50%-read/50%-write traffic. This is because the contemporary CPUs deploy the write-allocate cache policy [40]. With this policy, each store instruction first reads data from the main memory to the cache, then modifies it, and finally writes it to the main memory once the cache line is evicted. Each store instruction, therefore, does not correspond to a single memory write, but to one read and one write. [^1] Memory traffic with more than 50% of writes can be generated with non-temporal (streaming) stores that directly store data in the main memory. These instructions, however, are not supported in all HPC platforms. Even when supported, they are seldom used. In SPEC CPU2006 [41], SPEC CPU2017 [42], and SPLASH-2 [43] benchmarks running on the Intel Skylake server with Intel Pin tool [44], the non-temporal stores correspond to $<$1% of the overall memory instructions.
@@
The Mess bandwidth–latency curves are illustrated in Figure 1 (middle). The curves with different composition of read and write traffic are plotted with different shades of blue. Each curve is constructed based on tens of measurement points that cover the whole range of memory-traffic intensity. Figure 1 (top) illustrates the construction process for one of the curves. The memory access latency is measured with a pointer-chase benchmark [45, 46] executed on one CPU core or one GPU Stream Multiprocessor (SM). This determines the $y$-axis position of the measurement. The $x$-axis corresponds to memory bandwidth, monitored with hardware counters [47, 48, 49, 50]. Running pointer-chase alone measures the unloaded memory access latency. To measure the latency of the loaded memory system, concurrently with the pointer-chase, on the remaining CPU cores or GPU SMs, we run a memory traffic generator. The traffic generator is designed to create a wide range of memory traffic with configurable memory bandwidth utilization and read/write ratio. Both, pointer-chase and traffic generator are implemented in assembly to minimize any compiler intervention. The detailed implementations of pointer-chase and traffic generator is presented in Appendix -A. To minimize the latency penalties introduced by the TLB misses and the page walk, the Mess data-structures are allocated in huge memory pages. Additionally, at runtime, these overheads are monitored with hardware counters and subtracted from the memory latency measurements. The Mess benchmark release includes the source code and its detailed description, as well as the instructions for the benchmark compilation and execution.
@@ II-B Validation
The Mess benchmark provides a holistic and detailed memory characterization exceeding all existing benchmarks and tools. Still, these tools can validate some of the Mess measurements.
The unloaded memory system latency can be measured with LMbench and Google multichase in CPU platforms, and P-chase [51] in GPUs. We used these benchmarks to validate the Mess unloaded latency measurements in all hardware platforms under study, see Table I. In all experiments, Mess closely matches the LMbench, Google multichase and P-chase results.
@@
The Intel Memory Latency Checker (MLC) [11], can measure the memory latency for a selected memory traffic intensity, i.e. memory bandwidth. The memory bandwidth can be fine-tuned, but the tool provides a sparse analysis of different traffic compositions, i.e. read and write memory operations. We compare the Intel MLC results and the corresponding subset of the Mess measurements for all Intel platforms under study. The MLC and Mess results show the same trend, with slightly lower ($<$5%) latencies reported by the Mess benchmark. This is because the Mess is designed for close-to-the-hardware memory characterization, so it excludes the latency penalties introduced by the OS overheads, the TLB misses and the page walk.
@@ II-C Performance analysis
Figure 2 illustrates the Mess benchmark performance analysis with an example of the 24-core Intel Skylake server with six DDR4-2666 memory channels [52]. The figure confirms a general trend of the memory access latency which is initially roughly constant and then increases with higher memory pressure (i.e. bandwidth) due to resource contention among parallel accesses [36, 53, 54].
Detailed and close-to-the-hardware Mess characterization reveals some memory system aspects not discussed by previous studies. The most important one is the impact of the read and write memory traffic. The best performance, the lowest latency and the highest achieved bandwidth, are obtained for 100%-read traffic. Memory writes reduce the memory performance and reach the saturation point sooner. This is due to the extra timing constraints such as $t_{WR}$ and $t_{WTR}$, which come with memory write operations [55]. We detect this behavior for all the Intel, AMD, IBM, Fujitsu, and Amazon servers as well as NVIDIA GPUs used in the study with DDR4, DDR5, HBM2 and HBM2E memory (Section II-D). However, we see a very different write traffic impact on CXL memory expanders [33]. We analyze this behavior of CXL memory expanders in Section III-C.
Apart from the memory bandwidth–latency curves, we use the Mess benchmark to derive memory system performance metrics, also depicted in Figure 2, for quantitative comparison of different memory systems. In addition to the commonly-used unloaded memory latency, the detailed Mess characterization quantifies the maximum latency range of memory access latencies for all read-to-write ratios and the saturated bandwidth range. The memory system is saturated when any further increase in the memory system pressure, i.e. bandwidth, leads to a high increase in the memory access latency. We consider that the saturated bandwidth area starts at the point in which the memory access latency doubles the unloaded latency. In some of the platforms under study, we also detect a memory system over-saturation. Any increase in the memory system pressure after the over-saturation point causes a decline in the used memory bandwidth, while the access latency continues to increase. This behavior is observed in the bandwidth–latency curves as a “wave form” seen in Figure 2. The causes for this memory behavior are analyzed in Section II-D. To the best of our knowledge, our study is the first one to detect and analyze this suboptimal memory system behavior.
Although STREAM is the de facto standard for measuring sustained memory bandwidth [5, 56], Figure 2 shows significant gap between the STREAM bandwidth and the maximum measurements of the Mess benchmark. The sources of this gap are analyzed in the next section.
Figure 2: The Mess benchmark models the memory system performance with a family of bandwidth–latency curves. It covers the whole feasible read/write ratio of memory operations, and the whole space of the traffic intensity, from unloaded to fully-saturated memory systems.
@@
Stream kernels 53–61% 51–57% 46–51% 32–36% 78–82% 63–66% 49–55% 64–69%
@@
II-D Performance characterization: Actual systems
We use the Mess measurements to compare the memory system performance of Intel, AMD, IBM, Fujitsu and Amazon servers as well as NVIDIA GPUs with DDR4, DDR5, HBM2 and HBM2E. The platform and memory system characteristics are listed in Table I while Figure 3 shows their bandwidth–latency curves. The unloaded memory latency varies significantly between different platforms. It ranges from 85 ns in the Cascade Lake server with DDR4, to 122 ns in the A64FX servers with HBM2 and 129 ns in the Graviton 3 with DDR5 main memory. This difference should not be directly associated to the memory technology and standard. For example, the AMD Zen2 comprises DDR4 technology with practically the same command latencies (in nanoseconds) as the Intel Cascade Lake server, and still it shows almost 30 ns higher unloaded memory latency. This is because the load-to-use latency considers the time memory requests spend within the CPU chip, including the cache hierarchy and network on chip. These timings can differ significantly between different CPU architectures independently of the target memory system. Actually, we detect the highest unloaded memory latencies in the chips with the largest number of cores (Zen2, A64FX, Graviton 3) indicating that at least a portion of this latency is likely to be attributed to the network on chip which is larger and more complex in these architectures. NVIDIA H100 GPU shows higher unloaded latency due to its massive number of arithmetic processing units, lower on-chip frequency, and complex memory hierarchy [51, 63].

### v4

@@ Abstract
The current Mess benchmark release covers all major CPU and GPU ISAs, x86, ARM, Power, RISC-V, and NVIDIA’s PTX. We also release as open source the ZSim, gem5 and OpenPiton Metro-MPI integrated with the Mess simulator for DDR4, DDR5, Optane, HBM2, HBM2E and CXL memory expanders. The Mess application profiling is already integrated into a suite of production HPC performance analysis tools.
@@ I Introduction
Our study argues that the memory system benchmarking, simulation and application profiling can and should be based on a unified view of memory system performance. We provide this view with the Memory stress (Mess) framework comprised of the Mess benchmark, analytical memory system simulator and application profiling tool (Figure 1). Mess benchmark (Section II) describes the memory system performance with a family of bandwidth–latency curves. The benchmark covers the full range of the memory traffic intensity, from the unloaded to fully-saturated memory system. It also considers numerous compositions of read and write operations, plotted with different shades of blue in Figure 1 (middle). The Mess benchmark is designed for holistic and detailed memory system characterization, and it is easily adaptive to different target platforms. The current benchmark release covers all major CPU and GPU ISAs: x86, ARM, Power, RISC-V, and NVIDIA’s Parallel Thread Execution (PTX) [32].
@@
We deploy the Mess benchmark to characterize Intel, AMD, IBM, Fujitsu and Amazon servers as well as NVIDIA GPUs with DDR4, DDR5, HBM2 and HBM2E memory (Section III). We report and discuss a wide range of memory system behavior even for the hardware platforms with the same main memory configuration. These differences are especially pronounced in the high-bandwidth areas which have the greatest impact on memory-intensive applications. We also detect and analyze scenarios is which increase of the memory request rate leads to the degradation of the measured bandwidth.
@@
We use the Mess benchmark to evaluate memory system simulation of event-based ZSim [15], cycle-accurate gem5 [16] and RTL simulator OpenPiton Metro-MPI [33] (Section IV) We evaluate different internal memory models and widely-used external memory simulators, DRAMsim3 [22], Ramulator [22] and Ramulator 2 [24]. Unfortunately, all tested memory simulators, including well-established and trusted gem5 DDR models and cycle-accurate memory simulators, poorly resemble the actual system performance. The simulators show an unrealistically low load-to-use latency (starting at 4 ns), high memory bandwidth (exceeding 1.8$\times$ the maximum theoretical one), and a simulation error of tens of percents for memory-intensive benchmarks, STREAM [5], LMbench [8] and Google multichase [9]. We detect two sources of these errors: the Zsim interface towards the external memory simulators, and the imprecise DRAMsim3 and Ramulator model of the row-buffer utilization. Finally, although it was not the initial design target, we also detected a case in which holistic and detailed Mess benchmarking diagnosed a bug in the coherency protocol generated by the OpenPiton framework.
@@
Apart from the memory system characterization, the Mess bandwidth–latency curves can be also used for the memory performance simulation (Section V). We develop and integrate the Mess simulator in the ZSim, gem5, and OpenPiton Metro-MPI simulators, enabling simulation of high-end memory systems based on DDR4, DDR5, Optane, HBM2, and HBM2E technologies, and Compute Express Link (CXL) [34]. The Mess integration is easy, based on the standard interfaces between the CPU and external memory simulators [15, 16, 35]. The Mess simulator closely matches the actual memory systems performance. The simulation error for STREAM, LMbench and Google multichase benchmarks is between 0.4% and 6%, which is significantly better than any other memory models we tested. Mess is also fast. For example, Mess integrated with ZSim introduces only 26% simulation time increment over the fixed-latency memory model, while the speed-up over the Ramulator and DRAMsim3 ranges between 13$\times$ and 15$\times$. Mess also removes the current lag between the emergence of memory technologies and development of the reliable simulation models. For example, Mess is the first memory simulator that models CXL memory expanders, enabling further research on these novel memory devices. The simulation is based on the bandwidth–latency curves obtained from the memory manufacturer’s SystemC hardware model (Section V-C).
@@
Finally, the Mess memory bandwidth–latency curves enhance application profiling and performance analysis (Section VI). Mess application profiling determines positions of application execution time segments on the corresponding memory bandwidth–latency curves. The application memory stress can be combined with the overall application timeline analysis and can be linked to the source code. The Mess application profiling is already integrated into a suite of production HPC performance analysis tools [36].
@@
The inherent dependency between the used memory bandwidth and the access latency is by no means a new phenomenon, and the community has known about it at least for a couple of decades [37]. The Mess framework extends the previous work in three important aspects. Previous studies typically use a single bandwidth–latency memory curve to illustrate a general memory system behavior [37, 27, 12, 38, 39, 40]. The Mess benchmark is designed for holistic, detailed and close-to-the-hardware memory system performance characterization of the particular system under study. To reach this objective, the Mess benchmark is developed directly in the assembly, and the experiments are tailored to minimize and mitigate the impact of the system software. It detects and quantifies some aspects of the memory systems behavior not discussed in previous studies, such as the impact of the read and write memory traffic on performance, or discrepancies between different memory systems, both in actual platforms and hardware simulators. Second, the Mess framework tightly integrates the memory performance characterization into the memory simulation. The Mess simulator avoids complex memory system simulation and analytically adjusts the rate of memory instructions (provided by the CPU simulator) to the actual memory performance. The accuracy of this simulation approach relies on the input memory performance characterization. This characterization therefore has to be holistic, detailed and specific to the system under study, closely matching the Mess benchmark design. Third, the framework closely couples the memory-related profiling of hardware platforms and applications. Similar to the Mess simulator, the Mess application profiling itself is uncomplicated, and its real value comes from the application analysis in the context of the memory system characteristic. The quality of the memory characterization therefore directly impacts the the quality of the overall analysis. For this reason, the analysis has to be performed with detailed Mess-like memory performance description.
@@
The Mess benchmark is released as open source and it is ready to be used in x86, Power, ARM, and RISC-V CPUs and NVIDIA GPUs [32]. The release also contains all bandwidth–latency measurements shown in the paper, including the CXL expander curves provided by the memory manufacturer. We also release as open source the ZSim, gem5 and OpenPiton Metro-MPI integrated with the Mess simulator supporting DDR4, DDR5, Optane, HBM2, HBM2E and CXL expanders [41]. Also, the public releases of the production HPC performance analysis tools already include the Mess application profiling extension [42]. The released tools are ready to be used by the community for better understanding of the current and exploration of future memory systems.
@@ II-A Mess benchmark: Memory bandwidth–latency curves
The Mess memory characterization comprises tens of bandwidth–latency curves, each corresponding to a specific ratio between the read and write memory traffic. The Mess benchmark kernels cover the whole range of memory operations, from 100%-loads to 100%-stores. The 100%-load kernel generates a 100%-read memory traffic, while the 100%-store kernel creates 50%-read/50%-write traffic. This is because the contemporary CPUs deploy the write-allocate cache policy [43]. With this policy, each store instruction first reads data from the main memory to the cache, then modifies it, and finally writes it to the main memory once the cache line is evicted. Each store instruction, therefore, does not correspond to a single memory write, but to one read and one write. [^1] Memory traffic with more than 50% of writes can be generated with streaming stores that directly write data to the main memory. The public Mess repository already includes the benchmark with x86 streaming (non-temporal) stores. We are working on the equivalent benchmark version for ARM and Power CPUs (zero cache line fill instructions), and NVIDIA GPUs.
@@
The Mess bandwidth–latency curves are illustrated in Figure 1 (middle). The curves with different composition of read and write traffic are plotted with different shades of blue. Each curve is constructed based on tens of measurement points that cover the whole range of memory-traffic intensity. Figure 1 (top) illustrates the construction process for one of the curves. The $x$-axis of the chart corresponds to memory bandwidth, monitored with hardware counters [44, 45, 46, 47]. The memory access latency is measured with a pointer-chase benchmark [48, 49] executed on one CPU core or one GPU Stream Multiprocessor (SM). This determines the $y$-axis position of the measurement. The pointer-chase is selected because it is simple and portable to different actual hardware platforms and simulators. Optionally, memory latency could be measured with instruction-based sampling approach  [50] available in some current architectures [51, 52, 28].
Running pointer-chase alone measures the unloaded memory access latency. To measure the latency of the loaded memory system, concurrently with the pointer-chase, on the remaining CPU cores or GPU SMs, we run a memory traffic generator. The memory access latency is still measured only for the pointer-chase benchmark. The purpose of the traffic generator is to create memory accesses that will collide with the pointer-chase in the loaded memory system. The generator is designed to create memory traffic with configurable memory bandwidth utilization and read/write ratio. Each CPU core and GPU SM traverses two separate arrays, one with load and one with store operations. Therefore, the overall memory traffic is complex, determined by a sequential accesses within each array, but also by the interleaving between memory request from distinct arrays. The Mess benchmark covers a large range of row-buffer hit/empty/miss rates, e.g. between 35/43/22% and 84/13/3% in state-of-the-art Intel architectures (Section IV-D).
Both, pointer-chase and traffic generator are implemented in assembly to minimize any compiler intervention. The detailed implementations of pointer- chase and traffic generator is presented in Appendix -A. To minimize the latency penalties introduced by the TLB misses and the page walk, the Mess data-structures are allocated in huge memory pages. Additionally, at runtime, these overheads are monitored with hardware counters and subtracted from the memory latency measurements. The Mess benchmark release [32] includes the source code and its detailed description.
@@ II-B Validation
The Mess benchmark exceeds the coverage of all existing memory benchmarks and tools. Still, these tools can validate some of the Mess measurements.
The unloaded memory system latency can be measured with LMbench and Google multichase in CPU platforms, and P-chase [53] in GPUs. We used these benchmarks to validate the Mess unloaded latency measurements in all hardware platforms under study. In all experiments, Mess closely matches the LMbench, Google multichase and P-chase results.
@@
The Intel Memory Latency Checker (MLC) [11], can measure the memory latency for a selected memory traffic intensity, i.e. memory bandwidth. The memory bandwidth can be fine-tuned, but the tool provides a sparse analysis of different traffic compositions, i.e. read and write memory operations. We compare the Intel MLC results and the corresponding subset of the Mess measurements for all Intel platforms under study. The MLC and Mess results show the same trend, with slightly lower ($<$5%) latencies reported by the Mess benchmark. This is because the Mess is designed for close-to-the-hardware memory characterization, and, unlike the MLC, it excludes the latency penalties introduced by the OS overheads, the TLB misses and the page walk.
@@ II-C Performance analysis
Figure 2 illustrates the Mess benchmark performance analysis with an example of the 24-core Intel Skylake server with six DDR4-2666 memory channels [54]. The figure confirms a general trend of the memory access latency, which is initially roughly constant and then increases with higher memory pressure (i.e. bandwidth) due to resource contention among parallel accesses [37, 55, 56].
Detailed and close-to-the-hardware Mess characterization reveals some memory system aspects not discussed by previous studies. The most important one is the impact of the read and write memory traffic. The best performance, the lowest latency and the highest achieved bandwidth, are obtained for 100%-read traffic. Memory writes reduce the memory performance and reach the saturation point sooner. This is due to the extra timing constraints such as $t_{WR}$ and $t_{WTR}$, which come with memory write operations [57]. We detect this behavior for all the Intel, AMD, IBM, Fujitsu, and Amazon servers as well as NVIDIA GPUs used in the study with DDR4, DDR5, HBM2 and HBM2E memory (Section III). However, we see a very different write traffic impact on CXL memory expanders [34]. This behavior of CXL memory expanders is analyzed in Section V-C.
Apart from the memory bandwidth–latency curves, we use the Mess benchmark to derive memory system performance metrics, also depicted in Figure 2, for quantitative comparison of different memory systems. In addition to the commonly-used unloaded memory latency, the detailed Mess characterization quantifies the maximum latency range of memory access latencies for all read/write ratios and the saturated bandwidth range. The memory system is saturated when any further increase in the memory system pressure, i.e. bandwidth, leads to a high increase in the memory access latency. We consider that the saturated bandwidth area starts at the point in which the memory access latency doubles the unloaded latency. In some of the platforms under study, we also detect that the increase in the Mess memory access rate causes a memory bandwidth decline, while the access latency continues to increase. This behavior is observed in the bandwidth–latency curves as a “wave form” seen in Figure 2. The causes for this memory behavior are analyzed in Section III. To the best of our knowledge, our study is the first one to detect and analyze this suboptimal memory system behavior.
Figure 2 also depicts the bandwidth reported by the STREAM benchmark (vertical dashed lines). STREAM and Mess provide complementary memory bandwidth analysis. While STREAM is the de facto standard for measuring application-level sustained memory bandwidth [5, 58], the Mess benchmark provides the microarchitecture view and considers all memory traffic visible by hardware counters. Differences between these two approaches are analyzed in the next section.
Figure 2: The Mess benchmark models the memory system performance with a family of bandwidth–latency curves.
@@
STREAM kernels: Application view 53–61% 51–57% 46–51% 32–36% 78–82% 63–66% 49–55% 64–69%
@@
III Performance characterization: Actual systems
We use the Mess measurements to compare the memory system performance of Intel, AMD, IBM, Fujitsu and Amazon servers as well as NVIDIA GPUs with DDR4, DDR5, HBM2 and HBM2E. The platform and memory system characteristics are listed in Table I while Figure 3 shows their bandwidth–latency curves. The unloaded memory latency varies significantly between different platforms. It ranges from 85 ns in the Cascade Lake server with DDR4, to 122 ns in the A64FX servers with HBM2 and 129 ns in the Graviton 3 with DDR5 main memory. This difference should not be directly associated to the main memory. For example, the AMD Zen2 comprises DDR4 technology with practically the same command latencies (in nanoseconds) as the Intel Cascade Lake server, and still it shows almost 30 ns higher unloaded memory latency. This is because the load-to-use latency considers the time memory requests spend within the CPU chip, including the cache hierarchy and network on chip. These timings can differ significantly between different CPU architectures. We detect the highest unloaded memory latencies in the chips with the largest number of cores (Zen2, A64FX, Graviton 3) indicating that at least a portion of this latency is likely to be attributed to the network on chip which is larger and more complex in these architectures. NVIDIA H100 GPU shows higher unloaded latency due to its massive number of arithmetic processing units, lower on-chip frequency, and complex memory hierarchy [53, 65].